OpenStudy (anonymous):

Which expression is the simplified form of the square root 338 times w to the eighteenth power.? 13w3the square root 2 times w. 2w9the square root 13 times w. 2w3the square root 13 times w. 13w9the square root of 2.

OpenStudy (anonymous):

how many times does 2 go in to 18 ?

OpenStudy (anonymous):

i.e. what is \(18\div 2\) ?

OpenStudy (anonymous):

this is the easy part, the hard part is finding \(\sqrt{338}\)

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok so \(\sqrt{w^{18}}=w^9\) is a start

OpenStudy (anonymous):

as for \(\sqrt{338}\) it turns out that \(338=13\times 13\times 2\) or \(13^2\times 2\) that means \[\sqrt{338}=\sqrt{13^2\times 2}=\sqrt{13^2}\sqrt{2}=13\sqrt{2}\]

OpenStudy (anonymous):

giving you a final answer of \(13w^9\sqrt{2}\)

OpenStudy (anonymous):

which i hope is one of your choices

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then what ?

OpenStudy (anonymous):

looks like A is \(13w^3\sqrt{2}\) which is not right

OpenStudy (anonymous):

find the one that looks like \(13w^9\sqrt{2}\)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

okay pick that one

OpenStudy (anonymous):

thanks